Two trains on separate tracks move toward each other. Train 1 has a speed of 141 km/h; train 2, a speed of 85.0 km/h. Train 2 blows its horn, emitting a frequency of 500 Hz. What is the frequency heard by the engineer on train 1?
Hz
Answer #1

Doppler eqn states
First convert speeds to m/s so 141km/h = 141*(1000m/km)*(1h/3600s) = 39.166m/s

and 85.0km/h*(1000m/km)*(1h/3600s) = 23.6

fL = fS*(v + vL)/(v + vS) = 500Hz*(345 + 39.166)/(345-23.6) = 598.131Hz
Answer #2

velocity of sound = 330 m/s or 1188km/h

n' = (C+Vo)n/(C-Vs) = ( 1188+110) * 500 / ( 1188 - 70)
n' = 580.5 Hz.
Answer #3

velocity of sound = 330 m/s or 1188km/h

n' = (C+Vo)n/(C-Vs) = ( 1188+141) * 500 / ( 1188 - 85)
n' = 602.44Hz.